OpenStudy (anonymous):

If a varies directly as b, and a = 28 when b = 7, find b when a = 5.

OpenStudy (anonymous):

a = kb 28 = k*7 k = 28/7=4 a = 4*b plug in and see what you get

OpenStudy (imtant):

a= 28, b=7 a= 5 , b=? just multiply crossly. 28b=35 So b=1.25.